Road Warrior Corporation began operations early in the current year, building luxury motor homes. During the year, the company started and completed 50 motor homes at a cost of $60,000 per unit. Of these, 48 were sold for $100,000 each and two remain in finished goods inventory. In addition, the company had six partially completed units in its factory at year-end. Total costs for the year (summarized alphabetically) were as follows Direct materials used Direct labor General and administrative $ 721,000 841,000 500,000 expenses Income tax expense Manufacturing overhead Selling expenses 100,000 1,650,000 500,000 Instructions a. Compute the total manufacturing costs charged to work in process for the current year Total manufacturing costs b. Compute the cost of finished goods manufactured for the current year: Cost of finished goods manufactured c. Compute the cost of goods sold for the current year: Cost of goods sold

c. Compute the cost of goods sold for the current year: Cost of goods sold d. Compute the gross profit on sales for the current year:
Answer #1
a
Direct materials used 721000
Direct labor 841000
Manufacturing overhead 1650000
Total manufacturing costs 3212000
b
Cost of finished goods manufactured 3000000 =50*60000
c
Cost of goods sold 2880000 =48*60000
d
Sales revenue 4800000 =48*100000
Less: Cost of goods sold 2880000
Gross Profit 1920000